OpenStudy (anonymous):

Solve using any method. -x^2 + 2x - 2 = 0

OpenStudy (e.mccormick):

So what did you try?

OpenStudy (anonymous):

Nothing yet.

OpenStudy (e.mccormick):

Well, what method do you think you will need to use?

OpenStudy (anonymous):

since it's a quadratic function, what methods could you possibly use to solve for x?

OpenStudy (anonymous):

quadratic formula?

OpenStudy (e.mccormick):

With this, that could work well.

OpenStudy (anonymous):

okay now what?

OpenStudy (e.mccormick):

Well, know how to put it in that?

OpenStudy (anonymous):

determine values for a, b and c then substitute into the quadratics formula

OpenStudy (anonymous):

This is what I am unsure on.

OpenStudy (e.mccormick):

\(\dfrac{-b\pm \sqrt{b^2-4ac}}{2a}\)

OpenStudy (anonymous):

since it's -x^2 + 2x - 2 = 0 then; \[a=-1\] \[b=2\] \[c=-2\]

OpenStudy (anonymous):

substitute these values into the quadratics formula that e.mccormick has stated

OpenStudy (e.mccormick):

And are you up to complex (imaginary) numbers yet? The stuff with i?

OpenStudy (anonymous):

\[-2 \pm \sqrt{2^{2-}-4\times=1\times-2}\]

OpenStudy (anonymous):

divided by -2

OpenStudy (anonymous):

oops 2^2 - 4 x -1 x -2 *

OpenStudy (e.mccormick):

Ummm... OK. I see where the mistake there is. You hit = instad of -. \(\dfrac{-2\pm \sqrt{2^2-4(-1)(-2)}}{2(-1)}\)

OpenStudy (anonymous):

Yeah

OpenStudy (e.mccormick):

Now simplify it. =)

OpenStudy (anonymous):

ok

OpenStudy (e.mccormick):

The joy of algebra... and calculus too. You do one easy step. Then you spend the next several minutes simplifying what that did!

OpenStudy (anonymous):

\[-2\pm \sqrt{-4}\]

OpenStudy (anonymous):

over -2

OpenStudy (e.mccormick):

OK. And that simplifies more. Though that is a good start.

OpenStudy (anonymous):

\[(-2\pm2i)/-2\]

OpenStudy (e.mccormick):

Well, that looks like you can make it simpler still. Lot of 2s there.

OpenStudy (anonymous):

\[2i\]

OpenStudy (anonymous):

?

OpenStudy (e.mccormick):

Nope. But look at what you wrote this way: \(\dfrac{-2\pm 2i}{-2}\) What in that cancels?

OpenStudy (anonymous):

-2+2

OpenStudy (e.mccormick):

Is that after the cancel or just what cancels?

OpenStudy (anonymous):

Not sure...sorry

OpenStudy (e.mccormick):

\(\dfrac{-2\pm 2i}{-2}\implies \dfrac{-2}{-2}\cdot \dfrac{1\mp 1i}{1}\) That help?

OpenStudy (e.mccormick):

All I did was factor out the -2 on both the top and the bottom. On the top it makes the plus or minus become a minus or plus. That does not really matter because it is still both! But it lets you cancel out the -2 completely.

OpenStudy (anonymous):

Okay makes sense.

OpenStudy (e.mccormick):

So, have a final answer now?

OpenStudy (anonymous):

basically, the answer is \[1 \pm i \]

OpenStudy (anonymous):

or flip the plus/minus sign, it doesnt matter either way

OpenStudy (e.mccormick):

Yep. Here is how I worked it: \(\dfrac{-b\pm \sqrt{b^2-4ac}}{2a}\) \(\dfrac{-2\pm \sqrt{2^2-4(-1)(-2)}}{2(-1)}\) \(\dfrac{-2\pm \sqrt{4-8}}{-2}\) \(\dfrac{-2\pm \sqrt{-4}}{-2}\) \(\dfrac{-2\pm \sqrt{4}\sqrt{-1}}{-2}\) \(\dfrac{-2\pm 2\sqrt{-1}}{-2}\) \(\dfrac{2\mp 2\sqrt{-1}}{2}\) \(1\mp \sqrt{-1}\) \(1\mp i\) \(\therefore x=(1+i,1-i)\)

OpenStudy (anonymous):

Okay yeah! I get it thanks so much!

OpenStudy (e.mccormick):

The quadratic can be a pain in the anatomy, but it is good to remember because sometimes it is just your best choice. have fun!